# SIU3R: Simultaneous Scene Understanding and 3D Reconstruction Beyond Feature Alignment

### v1

@@ Abstract
Simultaneous understanding and 3D reconstruction plays an important role in developing end-to-end embodied intelligent systems. To achieve this, recent approaches resort to 2D-to-3D feature alignment paradigm, which leads to limited 3D understanding capability and potential semantic information loss. In light of this, we propose SIU3R, the first alignment-free framework for generalizable simultaneous understanding and 3D reconstruction from unposed images. Specifically, SIU3R bridges reconstruction and understanding tasks via pixel-aligned 3D representation, and unifies multiple understanding tasks into a set of unified learnable queries, enabling native 3D understanding without the need of alignment with 2D models. To encourage collaboration between the two tasks with shared representation, we further conduct in-depth analyses of their mutual benefits, and propose two lightweight modules to facilitate their interaction. Extensive experiments demonstrate that our method achieves state-of-the-art performance not only on the individual tasks of 3D reconstruction and understanding, but also on the task of simultaneous understanding and 3D reconstruction, highlighting the advantages of our alignment-free framework and the effectiveness of the mutual benefit designs.
@@ 3.1 Problem Formulation and Pipeline
SIU3R processes sparse unposed multi-view images with corresponding camera intrinsics , where is 2 in our setting and denotes the number of input context views, and learns a feed-forward network parameterized by network weights and learnable unified queries . The network establishes two key outputs: 1) pixel-aligned multi-view 3D Gaussians for 3D reconstruction, where is a single gaussian primitive and specify the image resolution; 2) mask prediction logits where , and class prediction logits where for scene understanding and indicates total classes including background. Formally, the learning objective implements the mapping:
@@ 3.3 Mutual Benefit Mechanism
Multi-View Mask Aggregation. Due to potential large viewpoint changes, occlusions, or lighting differences between images from different camera views, the semantic information may vary significantly. Thus, the masks directly predicted by the Unified Query Decoder may exhibit inconsistencies across different views. A typical case is shown in Fig.4 (a), the Unified Query Decoder only considers cross-view semantic correlation of 2D pixels, without realizing that these pixels from different views are probably neighbors in 3D space for the same instance. As a result, the same instance is predicted with different IDs across different views when semantic information alone is insufficient.
Benefiting from the pixel-aligned representation shared between our predicted 2D masks and 3D Gaussians, we can use 3D Gaussians as additional geometric clues to fuse multi-view 2D semantic information in 3D space and propagate it back to the original views to avoid inconsistency.
We call this “Reconstruction Helps Understanding (RU)”.
@@ 3.4 Training Objective
where and are rasterized and ground truth images, is derived from [55, 39]. In our training, we leverage both photometric loss and segmentation loss to simultaneously supervise 3D reconstruction and understanding. We set to 1, 0.5, 0.05, 0.05, 1, respectively.
@@ 4.1 Experimental Setup
Implementation Details. We utilize ScanNet[17] for training and validation, the largest public dataset that concurrently provides multi-view images with dense semantic/instance segmentation labels and text-referred segmentation labels[56]. We conduct training on 8 NVIDIA GeForce RTX 4090 GPUs, with our model trained for 100 epochs using a per-GPU batch size of 3 (total batch size of 24) for about 2 hours. In contrast, LSM[16] trains on 8 NVIDIA A100 GPUs for 3 days. AdamW optimizer[57] is employed with an initial learning rate of 1e-4 followed by cosine decay scheduling. All baseline methods are evaluated on ScanNet dataset under the same protocols as ours for fair comparison. More details can be found in our appendices in the supplementary material.
Baselines. Our experiments encompass not only isolated evaluation on separate tasks of 3D reconstruction and understanding, but also integrated evaluation on simultaneous understanding and 3D reconstruction. Therefore, we evaluate our method against three types of baseline methods, all of which are state-of-the-arts on their respective tasks: 1) Sparse-view 3D reconstruction: pixelSplat[29], MVSplat[30], NoPoSplat[37]; 2) Scene understanding: Mask2Former[39], LSeg[15]; 3) Simultaneous scene understanding and 3D reconstruction: LSM[16].
@@ 4.2 Main Results
The only method that can achieve 3D-aware understanding is LSM. However, its understanding capability is restricted by its source 2D model (LSeg) due to the nature of its feature alignment paradigm. Therefore, LSM can only support semantic and text-referred segmentation same as LSeg. Benefiting from our alignment-free paradigm and simultaneous task modeling, our generalist model supports both 2D and 3D understanding on comprehensive tasks including semantic, instance, panoptic, and text-referred segmentation within a unified framework, significantly exceeding other methods across all metrics.
Qualitative Results. The qualitative results also demonstrate the superiority of our method. As illustrated in Fig.5 (a), thanks to our Mask-Guided Geometry Refinement module that improves reconstructed 3D geometries, our method exhibits better visual quality and fewer artifacts than others in novel view synthesis. As demonstrated in Fig.5 (b), thanks to our simultaneous task modeling and Multi-View Mask Aggregation mechanism, our method can effectively leverage geometric clues to improve understanding, and thus outperform 2D-only methods like Mask2Former with less erroneous inclusions in semantic masks. Besides, since our alignment-free framework decouples us from off-the-shelf 2D vision-language models, the quality of our segmentation masks is much better than methods like LSM and LSeg that rely on language-based querying for segmentation. Similar effects can also be observed in text-referred segmentation results (Fig.5 (c)), where our methods significantly surpasses LSM and LSeg with sharper boundaries and less fragments. Please refer to our appendices for more results (3D instance and panoptic segmentation, extension to versatile 3D editing, comparisons with more 3D-based baselines, etc).
@@ 5 Conclusion
Limitations. Currently, SIU3R can only handle input of 2 views, which limits its application to global 3D understanding at larger scene scales. We plan to support video input with global scene observations as part of our future work. Additionally, since all the cross-view mask annotations in ScanNet dataset are obtained by projecting noisy 3D point clouds onto 2D images, the quality of our ground-truth labels is relatively poor, which may have a negative impact on our segmentation accuracy.
@@ Appendix B Comparisons with Per-scene Optimization Methods
We also compare our approaches to methods (i.e., Feature-3DGS[11] and NeRF-DFF[9]) that require dense view capturing and per-scene optimization. Both of the two per-scene optimization methods follow a feature alignment paradigm similar to the feed-forward method LSM[16], where their 3D understanding capabilities are powered by off-the-shelf 2D vision language models that can only support language-guided segmentation. To enable the training of Feature-3DGS and NeRF-DFF, we uniformly select dense views (i.e., 100 images) as input and conduct per-scene optimization for each scene to align 3DGS or NeRF field with 2D features via rasterization. As shown in Table II, our method surpasses all of the feature alignment-based approaches by a large margin in the task of scene understanding, no matter they perform reconstruction in per-scene optimization (Feature-3DGS and NeRF-DFF) or feed-forward (LSM) manner. Besides, benefiting from our align-free framework, our method can further enable instance-level understanding tasks such as instance and panoptic segmentation. Furthermore, our method is the fastest in reconstruction speed, significantly surpassing Feature-3DGS and NeRF-DFF, and leading ahead of LSM. Considering that Feature-3DGS and NeRF-DFF use much more training images than our method, our performance in novel view synthesis is acceptable while achieving the best depth accuracy owing to our mask-guided geometry refinement module. As shown in Fig.II, we also make qualitative comparisons with these feature alignment-based methods, where our method achieves superior mask quality and semantic coherence.
@@
Appendix C Additional Visualizations
C.1 Instance, Panoptic and Text-Referred Segmentation
@@
In our main manuscript, we have included qualitative comparisons with other methods and demonstrated our superiority in semantic segmentation. Here we provide additional qualitative results of instance and panoptic segmentation for further demonstration. As shown in Fig.III, compared to 2D-based method Mask2Former[39] that leads to noisy mask boundaries, our method exhibits significantly higher mask quality. As shown in Fig.III, when performing panoptic segmentation, our method exhibits excellent mask consistency across different views and significantly outperforms other methods in mask quality. Similar effects can also be observed in text-referred segmentation results as shown in Fig.V. We attribute the superiority of our method to the simultaneous modeling of scene understanding and 3D reconstruction, which effectively leverages 3D geometric clues to aggregate semantic information from different views, and propagates them back to the original views to ensure cross-view consistency and improve segmentation accuracy.
C.2 Depth Estimation
@@
As illustrated in Fig.VI, compared to other feed-forward reconstruction methods, our approach achieves significantly superior depth quality with less artifacts and better coherence. We attribute this to our mask-guided geometry refinement module, which ensures geometry consistency within the same object instances under semantic guidance of 2D masks, and thus reduces erroneous depth variations that typically observed in other approaches.
C.3 Versatile 3D Editing
@@
As shown in Fig.VII, we present a comprehensive set of versatile 3D editing results, demonstrating SIU3R’s potential for diverse 3D manipulation applications. Furthermore, this capability establishes an effective baseline that bridges geometric reconstruction, scene understanding and manipulation in 3D environments.

### v2

@@ Abstract
Simultaneous understanding and 3D reconstruction plays an important role in developing end-to-end embodied intelligent systems. To achieve this, recent approaches resort to 2D-to-3D feature alignment paradigm, which leads to limited 3D understanding capability and potential semantic information loss. In light of this, we propose SIU3R, the first alignment-free framework for generalizable simultaneous understanding and 3D reconstruction from unposed images. Specifically, SIU3R bridges reconstruction and understanding tasks via pixel-aligned 3D representation, and unifies multiple understanding (segmentation) tasks into a set of unified learnable queries, enabling native 3D understanding without the need of alignment with 2D models. To encourage collaboration between the two tasks with shared representation, we further conduct in-depth analyses of their mutual benefits, and propose two lightweight modules to facilitate their interaction. Extensive experiments demonstrate that our method achieves state-of-the-art performance not only on the individual tasks of 3D reconstruction and understanding, but also on the task of simultaneous understanding and 3D reconstruction, highlighting the advantages of our alignment-free framework and the effectiveness of the mutual benefit designs.
@@ 3.1 Problem Formulation and Pipeline
SIU3R processes sparse unposed multi-view images with corresponding camera intrinsics , where 2 in our setting and denotes the number of input context views, and learns a feed-forward network parameterized by network weights and learnable unified queries . The network establishes two key outputs: 1) pixel-aligned multi-view 3D Gaussians for 3D reconstruction, where is a single gaussian primitive and specify the image resolution; 2) mask prediction logits where , and class prediction logits where for scene understanding and indicates total classes including background. Formally, the learning objective implements the mapping:
@@ 3.3 Mutual Benefit Mechanism
Multi-View Mask Aggregation. Due to potential large viewpoint changes, occlusions, or lighting differences between images from different camera views, the semantic information may vary significantly. Thus, the masks directly predicted by the Unified Query Decoder may exhibit inconsistencies across different views. A typical case is shown in Fig.4 (a), the Unified Query Decoder only considers cross-view semantic correlation of 2D pixels, without realizing that these pixels from different views are probably neighbors in 3D space for the same instance.
As a result, the same instance is predicted with different IDs across different views when semantic information alone is insufficient.
Benefiting from the pixel-aligned representation shared between our predicted 2D masks and 3D Gaussians, we can use 3D Gaussians as additional geometric clues to fuse multi-view 2D semantic information in 3D space and propagate it back to the original views to avoid inconsistency. We call this “Reconstruction Helps Understanding (RU)”.
@@ 3.4 Training Objective
where and are rasterized and ground truth images, is derived from [55, 39], , where is binary cross-entropy loss, is dice loss and is classification loss. We follow [55, 39] set , and to 5.0, 5.0, 2.0. In our training, we leverage both photometric loss and segmentation loss to simultaneously supervise 3D reconstruction and understanding. We set to 1, 0.5, 0.05, 0.05, 1, respectively.
@@ 4.1 Experimental Setup
Implementation Details. We utilize ScanNet[17] for training and validation, the largest public dataset that concurrently provides multi-view images with dense semantic/instance segmentation labels and text-referred segmentation labels[56]. We conduct training on 8 NVIDIA GeForce RTX 4090 GPUs, with our model trained for 100 epochs using a per-GPU batch size of 3 (total batch size of 24) for about 2 hours. AdamW optimizer[57] is employed with an initial learning rate of 1e-4 followed by cosine decay scheduling.
Baselines. Our experiments encompass not only isolated evaluation on separate tasks of 3D reconstruction and understanding, but also integrated evaluation on simultaneous understanding and 3D reconstruction. Therefore, we evaluate our method against three types of baseline methods, all of which are state-of-the-arts on their respective tasks: 1) Sparse-view 3D reconstruction: pixelSplat[29], MVSplat[30], NoPoSplat[37]; 2) Scene understanding: Mask2Former[39], LSeg[15]; 3) Simultaneous scene understanding and 3D reconstruction: LSM[16]. All baseline methods are evaluated on ScanNet dataset under the same protocols as ours for fair comparison. To be more specific, for pixelSplat, MVSplat, NoPoSplat, Mask2Former, and LSM, we re-train their models following the training protocols of their official implementations using the processed ScanNet dataset same as ours. For reconstruction-only methods (i.e., pixelSplat, MVSplat, NoPoSplat), we only use their original rendering losses for supervision. For understanding-only method (i.e., Mask2Former), we only use its mask losses for supervision. For LSeg, since it already possesses general vision-language understanding capabilities, we directly adopt its pre-trained weights for evaluation.
@@ 4.2 Main Results
The only method that can achieve 3D-aware understanding is LSM. However, its understanding capability is restricted by its source 2D model (LSeg) due to the nature of its feature alignment paradigm. Therefore, LSM can only support semantic and text-referred segmentation same as LSeg. Benefiting from our alignment-free paradigm and simultaneous task modeling, our generalist model supports both 2D and 3D understanding on comprehensive tasks including semantic, instance, panoptic, and text-referred segmentation within a unified framework, significantly exceeding other methods across all metrics. We also conduct experiments to validate the generalizability of our method to more input views, unseen data domains and real-world scenarios. Please refer to our appendices for more results.
Qualitative Results. The qualitative results also demonstrate the superiority of our method. As illustrated in Fig.5 (a), thanks to our Mask-Guided Geometry Refinement module that improves reconstructed 3D geometries, our method exhibits better visual quality and fewer artifacts than others in novel view synthesis. As demonstrated in Fig.5 (b), thanks to our simultaneous task modeling and Multi-View Mask Aggregation mechanism, our method can effectively leverage geometric clues to improve understanding, and thus outperform 2D-only methods like Mask2Former with less erroneous inclusions in semantic masks. Besides, since our alignment-free framework decouples us from off-the-shelf 2D vision-language models, the quality of our segmentation masks is much better than methods like LSM and LSeg that rely on language-based querying for segmentation. Similar effects can also be observed in text-referred segmentation results (Fig.5 (c)), where our methods significantly surpasses LSM and LSeg with sharper boundaries and less fragments. Please refer to our appendices for more results (3D instance and panoptic segmentation, extension to versatile 3D editing, comparisons with more 3D-based baselines, real-world scenarios, etc).
@@ 5 Conclusion
Limitations. Currently, our SIU3R model is only trained on limited data compared to methods such as DUSt3R[34] and VGGT[58], which hinders its generalizability to broader visual domains. Additionally, since all the cross-view mask annotations in ScanNet dataset are obtained by projecting noisy 3D point clouds onto 2D images, the quality of our ground-truth labels is relatively poor, which may have a negative impact on our segmentation accuracy. We hope to address this in the future by introducing higher-quality labels or employing self-supervised methods less dependent on labeled data.
Acknowledgements. This work was supported in part by NSFC under Grant 62202389, in part by a grant from the Westlake University-Muyuan Joint Research Institute, in part by the Westlake Education Foundation, in part by the Headquarters Management Science and Technology Project of State Grid Corporation of China (No. 52090025001L-170-ZN), and in part by the Tianjin Science and Technology Plan Project (No. 24YDPYSN00150).
@@
In this appendix, we provide additional content to complement the main manuscript:
@@ Appendix B Comparisons with Per-Scene Optimization Methods
We also compare our approaches to methods (i.e., Feature-3DGS[11] and NeRF-DFF[9]) that require dense view capturing and per-scene optimization. Both of the two per-scene optimization methods follow a feature alignment paradigm similar to the feed-forward method LSM[16], where their 3D understanding capabilities are powered by off-the-shelf 2D vision language models that can only support language-guided segmentation. To enable the training of Feature-3DGS and NeRF-DFF, we uniformly select dense views (i.e., 100 images) as input and conduct per-scene optimization for each scene to align 3DGS or NeRF field with 2D features via rasterization. As shown in Table II, our method surpasses all of the feature alignment-based approaches by a large margin in the task of scene understanding, no matter they perform reconstruction in per-scene optimization (Feature-3DGS and NeRF-DFF) or feed-forward (LSM) manner. Besides, benefiting from our align-free framework, our method can further enable instance-level understanding tasks such as instance and panoptic segmentation. Furthermore, our method is the fastest in reconstruction speed, significantly surpassing Feature-3DGS and NeRF-DFF, and leading ahead of LSM. Considering that Feature-3DGS and NeRF-DFF use much more training images than our method, our performance in novel view synthesis is acceptable while achieving the best depth accuracy owing to our mask-guided geometry refinement module. As shown in Fig.III, we also make qualitative comparisons with these feature alignment-based methods, where our method achieves superior mask quality and semantic coherence.
@@
Appendix C Extend to Multi-View Inputs
Building upon the insights of some works on feed-forward multi-view reconstruction (e.g., NoPoSplat[37], VGGT[58]), we make minor modifications to our model to support more input views for simultaneous understanding and 3D reconstruction. Specifically, compared to the model with two views, where cross-attention is only performed between the tokens of the two views, for multiple views, we perform cross-attention between the tokens of each view and the concatenated tokens of all other views. This is then followed by our Gaussian decoder and Unified Query Decoder, which predict multi-view 3D Gaussians and masks, respectively. We re-trained four variant models with different number of input views (i.e., 2, 4, 6, 8 views). The quantitative results, shown in the table III, demonstrate that as the number of views increases, our model exhibits improved performance in not only reconstruction but also 3D-aware segmentation on novel views. We think this improvement is primarily due to the availability of additional view information, which reduces geometry uncertainty, and enhances the quality of 3D reconstruction. This, in turn, boosts 3D scene understanding performance, thanks to our mutual-benefit mechanism (i.e., R U).
| #Views | InferTime | InferVRAM | TrainVRAM | AbsRel | RMSE | PSNR | SSIM | LPIPS | mIoU | mAP | PQ |
| 2 | 0.067s | 4G | 24G | 0.074 | 0.208 | 25.96 | 0.822 | 0.184 | 0.592 | 0.271 | 0.650 |
| 4 | 0.097s | 5G | 38.5G | 0.072 | 0.205 | 26.28 | 0.822 | 0.179 | 0.611 | 0.283 | 0.687 |
| 6 | 0.140s | 6G | 56.5G | 0.071 | 0.203 | 26.45 | 0.830 | 0.175 | 0.617 | 0.287 | 0.693 |
| 8 | 0.185s | 7G | 77.5G | 0.071 | 0.201 | 26.65 | 0.838 | 0.170 | 0.620 | 0.289 | 0.697 |
Appendix D Comparison with Other Methods
Since DUSt3R, MAST3R, and VGGT are only designed to reconstruct scenes as 3D points rather than 3D Gaussians / masks, we note that they cannot be evaluated for the tasks of novel view synthesis and multi-task understanding. The comparison results are shown in the table below.
As for reconstruction performance, although DUSt3R, MAST3R and VGGT used much more training data (DUSt3R: 9 datasets, MASt3R: 14 datasets, VGGT: 17 datasets) and ground-truth depths for explicit geometry supervision, our method is only slightly inferior to them in terms of depth accuracy. We think such performance gap is reasonable because our method learns reconstruction solely through the supervision of novel view synthesis (NVS) task and was trained only on a single dataset. Besides, it is worth noting that our method exhibits the best depth accuracy among all the methods that support the novel view synthesis task (NVS), which focuses more on the rendering quality rather than merely on the geometry quality.
As for computational efficiency, our method exhibits faster inference speed compared to DUSt3R, MAST3R, and VGGT. We attribute this to the adoption of complex point post-processing for DUSt3R and MASt3R, and the significantly larger model size for VGGT (ours: 650M, VGGT: 1.1B).
| Method | Representation | Task | Runtime | AbsRel | RMSE |
| DUSt3R[34] | Point | Depth | 0.14s | 0.058 | 0.178 |
| MASt3R[35] | Point | Depth | 1.09s | 0.057 | 0.175 |
| VGGT[58] | Point | Depth | 0.17s | 0.052 | 0.174 |
| MVSplat[30] | 3DGS | Depth & NVS | 0.06s | 0.170 | 0.392 |
| NoPoSplat[37] | 3DGS | Depth & NVS | 0.06s | 0.094 | 0.243 |
| LSM[16] | 3DGS | Depth & NVS & Segm. | 0.19s | 0.075 | 0.219 |
| Ours | 3DGS | Depth & NVS & Segm. | 0.07s | 0.074 | 0.208 |
Appendix E Additional Visualizations
E.1 Real-World Captured Data
We collect some multi-view images in real-world scenarios using hand-held cellphones to further validate our generalization capability. As shown in Fig.II, we observe that our zero-shot performance in real-world scenes is surprisingly good in both reconstruction and segmentation.
E.2 Instance, Panoptic and Text-Referred Segmentation
@@
In our main manuscript, we have included qualitative comparisons with other methods and demonstrated our superiority in semantic segmentation. Here we provide additional qualitative results of instance and panoptic segmentation for further demonstration. As shown in Fig.IV, compared to 2D-based method Mask2Former[39] that leads to noisy mask boundaries, our method exhibits significantly higher mask quality. As shown in Fig.IV, when performing panoptic segmentation, our method exhibits excellent mask consistency across different views and significantly outperforms other methods in mask quality. Similar effects can also be observed in text-referred segmentation results as shown in Fig.VI. We attribute the superiority of our method to the simultaneous modeling of scene understanding and 3D reconstruction, which effectively leverages 3D geometric clues to aggregate semantic information from different views, and propagates them back to the original views to ensure cross-view consistency and improve segmentation accuracy.
E.3 Depth Estimation
@@
As illustrated in Fig.VII, compared to other feed-forward reconstruction methods, our approach achieves significantly superior depth quality with less artifacts and better coherence. We attribute this to our mask-guided geometry refinement module, which ensures geometry consistency within the same object instances under semantic guidance of 2D masks, and thus reduces erroneous depth variations that typically observed in other approaches.
E.4 Versatile 3D Editing
@@
As shown in Fig.VIII, we present a comprehensive set of versatile 3D editing results, demonstrating SIU3R’s potential for diverse 3D manipulation applications. Furthermore, this capability establishes an effective baseline that bridges geometric reconstruction, scene understanding and manipulation in 3D environments.